A random sample is drawn from a normally distributed population of known standard deviation 5. Construct a 99.8% confidence interval for the population mean based on the information given (not all of the information given need be used).

a. n = 16, 𝑥̅= 98, 𝑠 = 5.6

b. n = 9, 𝑥̅= 98, 𝑠 = 5.6


1
Expert's answer
2022-06-03T12:55:25-0400

1. The critical value for "\alpha = 0.002" "z_c = z_{1-\alpha/2} = 3.0902"

The corresponding confidence interval is computed as shown below:



"CI=(\bar{X}-z_c\times\dfrac{\sigma}{\sqrt{n}}, \bar{X}+z_c\times\dfrac{\sigma}{\sqrt{n}})""=(98-3.0902\times\dfrac{5}{\sqrt{16}},""98+3.0902\times\dfrac{5}{\sqrt{16}})""=(94.137,101.863)"

Therefore, based on the data provided, the 95 confidence interval for the population mean is "94.137 < \mu <101.863," which indicates that we are 99.8% confident that the true population mean "\mu" is contained by the interval "(94.137,101.863)."


2. The critical value for "\alpha = 0.002" "z_c = z_{1-\alpha/2} = 3.0902"

The corresponding confidence interval is computed as shown below:



"CI=(\bar{X}-z_c\times\dfrac{\sigma}{\sqrt{n}}, \bar{X}+z_c\times\dfrac{\sigma}{\sqrt{n}})""=(98-3.0902\times\dfrac{5}{\sqrt{9}},""98+3.0902\times\dfrac{5}{\sqrt{9}})""=(92.850,103.150)"

Therefore, based on the data provided, the 95 confidence interval for the population mean is "92.850 < \mu <103.150," which indicates that we are 99.8% confident that the true population mean "\mu" is contained by the interval "(92.85,103.150)."